a rectangle is 6 meters long, and its perimeter is 20 meters. what is its width?
Found 2 solutions by user_dude2008, Sphinx pinastri:
Answer by user_dude2008(1862)

2L+2W=P

2(6)+2W=20

12+2W=20

2W=20-12

2W=8

W=8/2

W=4

width is 4

Answer by Sphinx pinastri(17)

The total length of 2 adjacent sides is 20/2 = 10 m
One of them is 6m
So the other is 4m